Saved Help Save & Exit Submit 10 Check my work 2 Mason Company has two manufacturing departments-Machining and Assembly. The company considers all of its manufacturing overhead costs to be fixed costs. It provided the following estimates at the beginning of the year as well as the following information with respect to Jobs A and B Estimated Data Kanufacturing overhead Direct labor hours Machine hours points Machining Assembly Total $12,500,000 625,000 $13, 125,000 525,000 520,000 25,000 500,000 500,000 20,000 eBook Job A Direct labor hours Machine hours Machining Assembly Total 10 Print Job B Direct 1abor hours Machine hours Machining Assembly Total References 12 15 Required 1. If Mason Company uses a plantwide predetermined overhead rate with direct labor-hours as the allocation base, how much manufacturing overhead cost would be applied to Job A? Job B? (Round your answers to the nearest whole dollar amount.) 2. Assume that Mason Company uses departmental predetermined overhead rates. The Machining Department is allocated based on machine-hours and the Assembly Department is allocated based on direct labor-hours. How much manufacturing overhead cost would be applied to Job A? Job B? (Round your intermediate calculations and final answers to 2 decimal places.) 1. Manufacturing overhead applied Job A Manufacturing overhead applied Job B 2. Manufacturing overhead applied Job A Manufacturing overhead applied Job B
Answer #1

1. Allocation of manufacturing overhead in Job A and B

Per direct labour hour rate

Machinery Dept = $12500000/25000 = $ 500 per hour

Assembly Dept = $625000/500000 = $1.25

Job A manufacturing expense allocation.

Machinery Dept = $500*5 hours = $2500

Assembly Dept = $1.25*10 hours = $12.5

Total = $2512.5

Job B manufacturing expense allocation.

Machinery Dept = $500*4 hours = $2000

Assembly Dept = $1.25*5 hours = $6.25

Total = $2006..25

Question 2 Answer.

Per machine hour rate

Machinery Dept = $12500000/500000 = $ 25 per hour

Assembly Dept = $625000/20000 = $31.25 per hour

Job A manufacturing expense allocation.

Machinery Dept = $25*11 hours = $275

Assembly Dept = $1.25*10 hours = $12.5

Total = $287.5

Job B manufacturing expense allocation.

Machinery Dept = $25*12 hours = $300

Assembly Dept = $1.25*5 hours = $6.25

Total = $306..25